# Representations of a quantum-deformed Lorentz algebra, Clebsch-Gordan map, and Fenchel-Nielsen representation of quantum complex flat connections at level-$\bm{k}$

### v1

@@ 4.3 T-cycle trace operator
The operators $\bm{L},\tilde{\bm{L}}$ are generalizations of of the Dehn twist operator of quantum Teichmüller theory studied in [1, 3, 5]. The unitary transformation relating $(\bm{L}_{s},\tilde{\bm{L}}_{s})$ and $(\bm{L}_{t},\tilde{\bm{L}}_{t})$ can be seen as realizing the A-move of Moore-Seiberg groupoid on $\mathcal{H}$, as a generalization from the result from quantum Teichmüller theory in [6].
The following S-cycle and T-cycle trace operators and their tilded partners are also useful in the following discussion
$\bm{L}_{s}^{(aux)}=e^{-2{\bm{Y}}+L_{3}+L_{4}}+e^{2{\bm{Y}}-L_{3}-L_{4}}+e^{{\bm{X}}},$ (4.79)
$\bm{L}_{t}^{(aux)}=e^{-2{\bm{Y}}-L_{2}-L_{3}}+e^{2{\bm{Y}}+L_{2}+L_{3}}+e^{{\bm{X}}}.$ (4.80)
They are Fourier transforms of (4.42) and (4.69). The derivation of these operator does not rely on the assumption $m_{3}+m_{4}\in 2\mathbb{Z}$ or $m_{2}+m_{3}\in 2\mathbb{Z}$, in contrast to $\bm{L}$.
5 Eigenstates of the trace operators
The unbounded operator $\bm{L}$ is defined on the domain $\mathfrak{D}(\bm{L})\subset\mathcal{H}\simeq L^{2}(\mathbb{R})\otimes\mathbb{C}^{k}$. Any functions $f(\mu,m)$ in $\mathfrak{D}(\bm{L})$ should satisfy
$e^{\frac{4\pi}{k}b\mu}f(\mu,m)\in L^{2}(\mathbb{R}),\qquad e^{-\frac{4\pi}{k}b\mu}f(\mu,m)\in L^{2}(\mathbb{R}),\qquad f(\mu+ib,m)\in L^{2}(\mathbb{R}),\qquad\forall m\in\mathbb{Z}/k\mathbb{Z}.$
Recall that $\mathrm{Re}(b)>0$, the conditions are satisfied for $f(\mu,m)$ decaying as fast as $e^{\mp\frac{4\pi}{k}\left[\mathrm{Re}(b)+\epsilon\right]\mu}$ ($\epsilon>0$) when $\mu\to\pm\infty$ and holomorpic in the strip $\left\{\mu\in\mathbb{C}\mid\mathrm{Im}(\mu)\in\left[0,\mathrm{Re}(b)\right]\right\}$ satisfying $f(\mu+ib,m)\in L^{2}(\mathbb{R})$. The Hermite functions $e^{-\mu^{2}/2}H_{n}(\mu)$ , $n=1,\cdots,\infty$ satisfy the requirements and make a densed domain in $L^{2}(\mathbb{R})$. Therefore the domain $\mathfrak{D}(\bm{L})$is dense in $L^{2}(\mathbb{R})\otimes\mathbb{C}^{k}$. Similarly, the adjoint $\tilde{\bm{L}}=\bm{L}^{\dagger}$ is also densely defined on $L^{2}(\mathbb{R})\otimes\mathbb{C}^{k}$. ${\bm{L}}$ and its adjoint $\tilde{\bm{L}}$ are commutative.
We would like to solve the eigen-equations:
$\bm{L}\alpha_{r}(\mu,m)=\ell(r)\alpha_{r}(\mu,m),\qquad\tilde{\bm{L}}\alpha_{r}(\mu,m)=\ell(r)^{*}\alpha_{r}(\mu,m),\qquad\ell(r)\in\mathbb{C}.$
for some eigenstates $\alpha_{r}$. We can always make the following parametrization of any $\ell\in\mathbb{C}$,
$\ell=r+r^{-1},$
where $r\in\mathbb{C}^{\times}$ is determined by $\ell$ up to a “Weyl refection” $r\to r^{-1}$.
We consider the following ansatz
$\alpha_{r}(\mu,m)=\frac{\varphi\left(-{\bf q}^{-1}y^{2}r,-\tilde{{\bf q}}^{-1}\tilde{y}^{2}\tilde{r}\right)}{\varphi\left(-{\bf q}y^{-2}r,-\tilde{{\bf q}}\tilde{y}^{-2}\tilde{r}\right)}f_{r}(\mu,m),$ (5.1)
where $r,\tilde{r}$ are parametrized as the following
$r=\exp\left[\frac{2\pi i}{k}(-ib\mu_{r}-m_{r})\right],\qquad\tilde{r}=\exp\left[\frac{2\pi i}{k}(-ib^{-1}\mu_{r}+m_{r})\right]\qquad\mu_{r}\in\mathbb{R},\quad m_{r}\in\mathbb{Z}/k\mathbb{Z}.$ (5.2)
The eigen-equations implies the following recursion relations of $f_{r}(\mu,m)$:
$rf_{r}(\mu+ib,m-1)=f_{r}(\mu,m),\qquad\tilde{r}f_{r}(\mu+ib^{-1},m+1)=f_{r}(\mu,m),$ (5.3)
and we find that
$f_{r}(\mu,m)=e^{-\frac{2\pi i}{k}\left(-\mu_{r}\mu+m_{r}m\right)},$ (5.4)
is a solution for any $r$. Note that in the above, we have restrict $r$ to satisfy $m_{r}\in\mathbb{Z}/k\mathbb{Z}$ so that $f_{r}(\mu,m)$ is single-valued. As a result, up to a constant rescaling independent of $\mu,m$, the following function
$\alpha_{r}(\mu,m)=e^{-\frac{2\pi i}{k}\left(-\mu_{r}\mu+m_{r}m\right)}\frac{\varphi\left(-{\bf q}^{-1}y^{2}r,-\tilde{{\bf q}}^{-1}\tilde{y}^{2}\tilde{r}\right)}{\varphi\left(-{\bf q}y^{-2}r,-\tilde{{\bf q}}\tilde{y}^{-2}\tilde{r}\right)}$ (5.5)
is the eigenstate of $\bm{L}$ and $\tilde{\bm{L}}$ corresponding to the eigenvalues $r+r^{-1}$ and $\tilde{r}+\tilde{r}^{-1}$.
A similar computation shows that the eignstates of $\bm{L}^{(aux)}_{s},\tilde{\bm{L}}^{(aux)}_{s}$ and $\bm{L}^{(aux)}_{t},\tilde{\bm{L}}^{(aux)}_{t}$ are given respectively by
$\alpha_{r}^{(s)}(\mu,m)$ $=$ $e^{-\frac{2\pi i}{k}\left(-\mu_{r}\mu+m_{r}m\right)}\frac{\varphi\left(-{\bf q}^{-1}y^{2}\lambda_{3}^{-1}\lambda_{4}^{-1}r,-\tilde{{\bf q}}^{-1}\tilde{y}^{2}\tilde{\lambda}_{3}^{-1}\tilde{\lambda}_{4}^{-1}\tilde{r}\right)}{\varphi\left(-{\bf q}y^{-2}\lambda_{3}\lambda_{4}r,-\tilde{{\bf q}}\tilde{y}^{-2}\tilde{\lambda}_{3}\tilde{\lambda}_{4}\tilde{r}\right)}\ ,$ (5.6)
$\alpha_{r}^{(t)}(\mu,m)$ $=$ $e^{-\frac{2\pi i}{k}\left(-\mu_{r}\mu+m_{r}m\right)}\frac{\varphi\left(-{\bf q}^{-1}y^{2}\lambda_{2}\lambda_{3}r,-\tilde{{\bf q}}^{-1}\tilde{y}^{2}\tilde{\lambda}_{2}\tilde{\lambda}_{3}\tilde{r}\right)}{\varphi\left(-{\bf q}y^{-2}\lambda_{2}^{-1}\lambda_{3}^{-1}r,-\tilde{{\bf q}}\tilde{y}^{-2}\tilde{\lambda}_{2}^{-1}\tilde{\lambda}_{3}^{-1}\tilde{r}\right)}\ .$ (5.7)
Theorem 5.1.
For $k=2N$ with odd $N$, the spectral decomposition of $(\bm{L},\tilde{\bm{L}})$, $(\bm{L}^{(aux)}_{s},\tilde{\bm{L}}^{(aux)}_{s})$, or $(\bm{L}^{(aux)}_{t},\tilde{\bm{L}}^{(aux)}_{t})$ gives the following direct integral representation
$\mathcal{H}\simeq\sum_{m_{r}\in\mathbb{Z}/k\mathbb{Z}}\int_{0}^{\infty}\mathrm{d}\mu_{r}\,\rho(\mu_{r},m_{r})^{-1}\,\mathcal{H}_{\mu_{r},m_{r}}.$ (5.8)
where each $\mathcal{H}_{\mu_{r},m_{r}}$ is 1-dimensional, and
$\rho(\mu_{r},m_{r})$ $=$ $N^{2}\left[\sin\left(\frac{\pi}{N}(ib\mu_{r}+m_{r})\right)\sin\left(\frac{\pi}{N}(-ib^{-1}\mu_{r}+m_{r})\right)\right]^{-1}$ (5.9)
Any $f\in\mathcal{H}$ can be represented by $f(\mu_{r},m_{r})$ where $\mu_{r}\in\mathbb{R}_{+}$ and $m_{r}\in\mathbb{Z}/k\mathbb{Z}$, such that $\bm{L},\tilde{\bm{L}}$ are represented as the multiplication operators
The same holds for $(\bm{L}^{(aux)}_{s},\tilde{\bm{L}}^{(aux)}_{s})$ or $(\bm{L}^{(aux)}_{t},\tilde{\bm{L}}^{(aux)}_{t})$. For any pair of states $f,f^{\prime}\in\mathcal{H}$
$\langle f\mid f^{\prime}\rangle=\sum_{m_{r}\in\mathbb{Z}/k\mathbb{Z}}\int_{0}^{\infty}\mathrm{d}\mu_{r}\,\rho(\mu_{r},m_{r})^{-1}f(\mu_{r},m_{r})^{*}f^{\prime}(\mu_{r},m_{r}).$ (5.11)
We call this direction integral representation the FN representation of $S$-cycle or $T$-cycle, since the corresponding trace operator is diagonalized.
For $k=2N$ with odd $N$, we have the decomposition $\mathcal{H}=\mathcal{H}_{+}\oplus\mathcal{H}_{-}$, where $f_{\pm}(\mu,m+N)=\pm f_{\pm}(\mu,m)$ for $f_{\pm}\in\mathcal{H}_{\pm}$. On each of $\mathcal{H}_{\pm}$, $\bm{x},\bm{y}^{2}$ are represented as
$\bm{y}^{2}f_{\pm}(\mu,m)=e^{\frac{2\pi i}{N}(-ib\mu-m)}f_{\pm}(\mu,m),$ $\quad\bm{x}f_{\pm}(\mu,m)=f_{\pm}(\mu+ib,m-1),$
$\tilde{\bm{y}}^{2}f_{\pm}(\mu,m)=e^{\frac{2\pi i}{N}(-ib^{-1}\mu+m)}f_{\pm}(\mu,m),$ $\quad\tilde{\bm{x}}f_{\pm}(\mu,m)=f_{\pm}(\mu+ib^{-1},m+1)$ (5.12)
where $f_{\pm}\in\mathcal{H}_{\pm}$ respectively and $m\in\mathbb{Z}/N\mathbb{Z}$. $\bm{L},\tilde{\bm{L}}$ leave each of $\mathcal{H}_{+},\mathcal{H}_{-}$ invariant. Correspondingly, the eigenstate $\alpha_{r}$ satisfy
$\alpha_{r}(\mu,m+N)=(-1)^{m_{r}}\alpha_{r}(\mu,m),$ (5.13)
i.e. it is $N$-periodic as $f_{+}$ (anti-periodic as $f_{-}$) when $m_{r}$ is even (odd).
$\mathcal{H}_{\pm}$ as subspaces of $\mathcal{H}$ have the inner products inherited from $\mathcal{H}$. Since ±$\mathcal{H}_{\pm}\simeq L^{2}(\mathbb{R})\otimes\mathbb{C}^{N}$ It is also helpful to consider a different inner product on $\mathcal{H}_{\pm}$:
$\langle f_{\pm}\mid f^{\prime}_{\pm}\rangle_{N}=\sum_{m\in\mathbb{Z}/N\mathbb{Z}}\int\mathrm{d}\mu\,f_{\pm}(\mu,m)^{*}f^{\prime}_{\pm}(\mu,m)=\frac{1}{2}\langle f_{\pm}\mid f^{\prime}_{\pm}\rangle.$ (5.14)
Since all $\bm{L},\tilde{\bm{L}}$, $(\bm{L}^{(aux)}_{s},\tilde{\bm{L}}^{(aux)}_{s})$, and $(\bm{L}^{(aux)}_{t},\tilde{\bm{L}}^{(aux)}_{t})$ leave each of $\mathcal{H}_{+},\mathcal{H}_{-}$ invariant, we prove Theorem 5.1 by studying the spectral decompostion in $\mathcal{H}_{+}$ and $\mathcal{H}_{-}$ separately. The following proof mainly focuses on the spectral decomposition of $\bm{L},\tilde{\bm{L}}$. The proof for $\bm{L}^{(aux)}_{s},\tilde{\bm{L}}^{(aux)}_{s}$ and $\bm{L}^{(aux)}_{t},\tilde{\bm{L}}^{(aux)}_{t}$ can be obtained analogously.
5.1 Periodic states
We firstly restrict to the subset of eigenstates $\alpha_{r}$ with even $m_{r}$. In this case, we can express $\alpha_{r}$ in terms of $\gamma(x,n)={\rm D}_{b}(-x,n)$, where ${\rm D}_{b}$ is the quantum dilogarithm function studied in [11, 10] (see Appendix A for some useful properties):
@@ 5.1 Periodic states
$\gamma(-x,n)$ $=$ ${\rm D}_{b}(x,n)=\prod_{j=0}^{\infty}\frac{1-{\bf q}^{2j+1}\exp\left[\frac{2\pi i}{N}\left(-ibx\sqrt{N}+n\right)\right]}{1-\tilde{{\bf q}}^{-2j-1}\exp\left[\frac{2\pi i}{N}\left(-ib^{-1}x\sqrt{N}-n\right)\right]},$ (5.15)
${\bf q}$ $=$ $e^{\frac{\pi i}{N}(1+b^{2})},\qquad\tilde{{\bf q}}=e^{\frac{\pi i}{N}(1+b^{-2})},$ (5.16)
and $\mathrm{D}_{b}$ relates to $\varphi$ by
@@
$\varphi\left(-y^{2},-\tilde{y}^{2}\right)=\mathrm{D}_{b}\left(\frac{\mu}{\sqrt{N}},-m\right)^{-1}=\gamma\left(-\frac{\mu}{\sqrt{N}},-m\right)^{-1}.$ (5.17)
The expression of $\alpha_{r}$ is given by
$\alpha_{r}(\mu,m)=e^{-\frac{\pi i}{N}\left[m_{r}m-\mu_{r}\mu\right]}\frac{\gamma\left(-\frac{\mu_{r}/2-\mu+c_{b}}{\sqrt{N}},m-m_{r}/2\right)}{\gamma\left(-\frac{\mu_{r}/2+\mu-c_{b}}{\sqrt{N}},-m-m_{r}/2\right)},\qquad c_{b}=\frac{i}{2}(b+b^{-1}).$ (5.18)
$\alpha_{r}$ has poles along the real $\mu$ axis, so we propose a regularization by shifting $\mu\to\mu+i\epsilon\sqrt{N}$ with $\epsilon>0$ inside $\gamma$. Moreover, if we multiply $\alpha_{r}$ by a factor $\beta_{r}$ independent of $\mu,m$
$\psi_{r}=\beta_{r}\alpha_{r},\qquad\beta_{r}=\exp\left(-\frac{i\pi c_{b}^{2}}{3N}-\frac{i\pi c_{b}\mu_{r}}{N}+\frac{i\pi m_{r}}{2}+\frac{i\pi\mu_{r}^{2}}{4N}-\frac{i\pi N}{6}\right),$ (5.19)
$\psi_{r}$ is manifestly invariant under the Weyl reflection $r\to r^{-1}$ (recall that the eigenvalue is invariant under the reflection ):
$\psi_{r}=\psi_{r^{-1}}.$ (5.20)
@@
Indeed, we introduce some short-hand notations  [^6] These notations are inspired by [7].
$\omega=\frac{i}{2b\sqrt{N}},\quad\omega^{\prime}=\frac{ib}{2\sqrt{N}},\quad\omega^{\prime\prime}=\frac{c_{b}}{\sqrt{N}},\quad x=\frac{\mu}{\sqrt{N}},\quad\lambda=\frac{\mu_{r}}{2\sqrt{N}}.$ (5.21)
and apply the inverse relation $\gamma(x,n)\gamma(-x,-n)=\exp(-\frac{i\pi n^{2}}{N}-i\pi n-\frac{i\pi N}{6}+i\pi x^{2}-\frac{1}{3}i\pi\omega^{\prime\prime}{}^{2})$ to $\alpha_{r}$, and we obtain
$\psi_{r}(\mu,m)$ $=$ $e^{\frac{i\pi m^{2}}{N}+\frac{i\pi m_{r}^{2}}{4N}-i\pi m-i\pi\left(x+i\epsilon-\omega^{\prime\prime}\right)^{2}}$ (5.22)
$\gamma\left(-\lambda+x-\omega^{\prime\prime}+i\epsilon,m-\frac{m_{r}}{2}\right)\gamma\left(\lambda+x-\omega^{\prime\prime}+i\epsilon,m+\frac{m_{r}}{2}\right)$
which is manifestly invariant under $(\lambda,m_{r})\to(-\lambda,-m_{r})$  [^7] A similar computation re-scales $\alpha_{r}^{(s)},\alpha_{r}^{(t)}$ to $\psi_{r}^{(s)},\psi_{r}^{(t)}$, where $\psi_{r}^{(s)}$ $=$ $e^{\frac{i\pi m_{r}^{2}}{4N}+\frac{i\pi\left(m-\frac{m_{3}}{2}-\frac{m_{4}}{2}\right){}^{2}}{N}-i\pi\left(m-\frac{m_{3}}{2}-\frac{m_{4}}{2}\right)-i\pi\left(x-\omega^{\prime\prime}+i\epsilon\right)^{2}}$ (5.23) $\gamma\left(-\lambda+x-\omega^{\prime\prime}+i\epsilon,-\frac{m_{r}}{2}+m-\frac{m_{3}}{2}-\frac{m_{4}}{2}\right)\gamma\left(\lambda+x-\omega^{\prime\prime}+i\epsilon,\frac{m_{r}}{2}+m-\frac{m_{3}}{2}-\frac{m_{4}}{2}\right)$ up to a constant shift of $x$, and $\psi_{r}^{(t)}$ is given by $m_{3}\to-m_{2},\ m_{4}\to-m_{3}$. Both $\psi_{r}^{(s)}$ and $\psi_{r}^{(t)}$ are invariant under $(\lambda,m_{r})\to(-\lambda,-m_{r})$.. We often fix the symmetry by requiring $\lambda\geq 0$ ($\mu_{r}\geq 0$).
@@
The following two lemmas state that $\psi_{r}$ for even $m_{r}$ form a complete distributional orthogonal basis for $\mathcal{H}_{+}$.
@@
The eigenstates with even $m_{r}$ satisfy the following orthogonality
$\left\langle\psi_{r}\mid\psi_{r^{\prime}}\right\rangle_{N}=\frac{1}{2}\left\langle\psi_{r}\mid\psi_{r^{\prime}}\right\rangle=\frac{1}{2}\rho(\mu_{r},m_{r})\delta_{r,r^{\prime}}.$ (5.24)
@@
$\delta_{r,r^{\prime}}$ $=$ $\frac{1}{2}\left[\delta(\mu_{r}-\mu_{r^{\prime}})\delta_{m_{r},m_{r^{\prime}}}+\delta(\mu_{r}+\mu_{r^{\prime}})\delta_{m_{r},-m_{r^{\prime}}}\right].$ (5.25)
@@
First of all, by the periodicity of $\psi_{r}$, we have
$\sum_{m\in\mathbb{Z}/k\mathbb{Z}}\int\mathrm{d}\mu\,\psi_{r}(\mu,m)^{*}\psi_{r^{\prime}}(\mu,m)=2\sum_{m\in\mathbb{Z}/N\mathbb{Z}}\int\mathrm{d}\mu\,\psi_{r}(\mu,m)^{*}\psi_{r^{\prime}}(\mu,m).$ (5.26)
@@
$e^{-\frac{i\pi(m_{r}^{2}-m_{r^{\prime}}^{2})}{4N}+4i\pi x(\omega^{\prime\prime}-i\epsilon)}\frac{\gamma\left(-\lambda^{\prime}+x-\omega^{\prime\prime}+i\epsilon,m-\frac{m_{r^{\prime}}}{2}\right)\gamma\left(\lambda^{\prime}+x-\omega^{\prime\prime}+i\epsilon,m+\frac{m_{r^{\prime}}}{2}\right)}{\gamma\left(-\lambda+x+\omega^{\prime\prime}-i\epsilon,m-\frac{m_{r}}{2}\right)\gamma\left(\lambda+x+\omega^{\prime\prime}-i\epsilon,\frac{m_{r}}{2}+m\right)},$ (5.27)
where $\lambda^{\prime}={\mu_{r^{\prime}}}/{2\sqrt{N}}$. We use the integration identity (A.17) to transform the ratio of two $\gamma$’s  [^8] The variables involved in (A.17) are given by $t=x$, $\alpha=\lambda^{\prime}-\omega^{\prime\prime}+i\epsilon,\ \beta=\lambda^{\prime}+\omega^{\prime\prime}-i\epsilon$, $d=m$, $a=m_{r}^{\prime}/2$, and $p=m_{r}/2$. We have $\mathrm{Im}\left(\alpha+\frac{c_{b}}{\sqrt{N}}\right)=\epsilon>0,\ \mathrm{Im}\left(\frac{c_{b}}{\sqrt{N}}-\beta\right)=\epsilon>0,\ \mathrm{Im}(\alpha-\beta)=2\epsilon-\frac{b+b^{-1}}{\sqrt{N}}<0$ for sufficiently small $\epsilon>0$. Therefore there exist an integration contour of $s$ with $\mathrm{Im}(\alpha-\beta)<\mathrm{Im}(s)<0$ allowing (A.17) to hold.
$\frac{\gamma\left(\lambda^{\prime}+x-\omega^{\prime\prime}+i\epsilon,m+\frac{m_{r^{\prime}}}{2}\right)}{\gamma\left(\lambda+x+\omega^{\prime\prime}-i\epsilon,\frac{m_{r}}{2}+m\right)}$ (5.28)
$=$ $\frac{\zeta_{0}}{\sqrt{N}}\sum_{c\in\mathbb{Z}/N\mathbb{Z}}\int\mathrm{d}s\frac{e^{\frac{2i\pi c(m+m_{r^{\prime}}/2)}{N}+2i\pi s(x+\lambda^{\prime})}\gamma\left(-\lambda^{\prime}+\lambda+s+\omega^{\prime\prime}-2i\epsilon,-c-\frac{m_{r^{\prime}}}{2}+\frac{m_{r}}{2}\right)}{\gamma\left(s+\omega^{\prime\prime},-c\right)\gamma\left(-\lambda^{\prime}+\lambda+\omega^{\prime\prime}-2i\epsilon,\frac{m_{r}}{2}-\frac{m_{r^{\prime}}}{2}\right)},$

### v2

@@ 4.4 T-cycle trace operator
The operators $\bm{L},\tilde{\bm{L}}$ are generalizations of of the Dehn twist operator of quantum Teichmüller theory studied in [1, 3, 5]. The unitary transformation $\mathfrak{U}_{s}^{-1}\mathfrak{U}_{t}$ relating $(\bm{L}_{s},\tilde{\bm{L}}_{s})$ and $(\bm{L}_{t},\tilde{\bm{L}}_{t})$ can be seen as realizing the A-move of Moore-Seiberg groupoid on $\mathcal{H}$, as a generalization from the result from quantum Teichmüller theory in [6].
5 Spectral decomposition
The above discuss shows that both the S-cycle and T-cycle trace operators of quantized flat connection on 4-holed sphere, $(\bm{L}_{s},\tilde{\bm{L}}_{s})$ and $(\bm{L}_{t},\tilde{\bm{L}}_{t})$ are unitary equivalent to $(\bm{L},\tilde{\bm{L}})$, which are derived from the unitary transformations acting on the Casimir operators $(Q_{21},\tilde{Q}_{21})$ of the tensor product representation of $\mathscr{U}_{\mathbf{q}}(sl_{2})\otimes\mathscr{U}_{\widetilde{\mathbf{q}}}(sl_{2})$ on $\mathcal{H}\otimes\mathcal{H}$. On one hand, the spectral decomposition of $(\bm{L},\tilde{\bm{L}})$ is unitary equivalent to the Clebsch-Gordan decomposition for $\mathscr{U}_{\mathbf{q}}(sl_{2})\otimes\mathscr{U}_{\widetilde{\mathbf{q}}}(sl_{2})$ by Lemma 3.3, while on the other hand, diagonalizing $(\bm{L},\tilde{\bm{L}})$ gives the Fenchel-Nielsen (FN) representation of $\mathcal{H}$, since classically $(L_{s},\tilde{L}_{s})$ and $(L_{t},\tilde{L}_{t})$ are the complexification of the Fenchel-Nielsen (FN) lengths on the 4-holed sphere, and $(\bm{L}_{s},\tilde{\bm{L}}_{s})$ and $(\bm{L}_{t},\tilde{\bm{L}}_{t})$ define their quantizations.
We define
@@ 5 Spectral decomposition
$\gamma(-x,n)$ $=$ $\prod_{j=0}^{\infty}\frac{1-{\mathbf{q}}^{2j+1}\exp\left[\frac{2\pi i}{N}\left(-ibx\sqrt{N}+n\right)\right]}{1-\widetilde{{\mathbf{q}}}^{-2j-1}\exp\left[\frac{2\pi i}{N}\left(-ib^{-1}x\sqrt{N}-n\right)\right]},$ (5.1)
$\gamma(-x,n)=\mathrm{D}_{b}(x,n)$ is the quantum dilogarithm over ×$\mathbb{R}\times\mathbb{Z}/N\mathbb{Z}$ defined by Andersen and Kashaev in [10] (see also [11]). For $N=1$, $\gamma(x,n)=\gamma(x,0)\equiv\gamma(x)$ is Faddeev’s quantum dilogarithm in e.g. [7, 27]:
$\gamma(x)=\exp\left[\frac{1}{4}\int_{\mathbb{R}+i0^{+}}\frac{\mathrm{d}w}{w}\frac{e^{2iwx}}{\sinh\left(b^{-1}{w}\right)\sinh(bw)}\right]$ (5.2)
The relation with the quantum dilogarithm $\varphi$ used above is
@@
$\varphi\left(-y,-\tilde{y}\right)=\gamma\left(-\frac{\mu}{\sqrt{N}},-m\right)^{-1},\qquad y=e^{\frac{2\pi i}{N}(-ib\mu-m)},\quad\tilde{y}=e^{\frac{2\pi i}{N}(-ib^{-1}\mu+m)}$ (5.3)
@@
We introduce some short-hand notations that are useful below
$c_{b}=\frac{i}{2}(b+b^{-1}),\quad\omega=\frac{i}{2b\sqrt{N}},\quad\omega^{\prime}=\frac{ib}{2\sqrt{N}},\quad\omega^{\prime\prime}=\frac{c_{b}}{\sqrt{N}},\quad x=\frac{\mu}{\sqrt{N}},\quad\lambda=\frac{\mu_{r}}{\sqrt{N}}.$ (5.4)
Some useful properties of $\gamma(x,n)$ including the inverse relation, recursion relation, and integral identity can be found in Appendix A.
$\bm{L},\widetilde{\bm{L}}$ has the following eigenstates
$\psi_{r}(\mu,m)$ $=$ $\exp\left(\frac{i\pi m^{2}}{N}+\frac{i\pi m_{r}^{2}}{N}+i\pi m_{r}^{2}-i\pi m-i\pi\left(x-\omega^{\prime\prime}+i\epsilon\right)^{2}\right)$ (5.5)
$\gamma\left(-\lambda+x-\omega^{\prime\prime}+i\epsilon,m-m_{r}\right)\gamma\left(\lambda+x-\omega^{\prime\prime}+i\epsilon,m_{r}+m\right),$
It is straight-forward to check that $\psi_{r}$ satisfies the following eigen-equations
$\bm{L}\psi_{r}(\mu,m)=\left(r+r^{-1}\right)\psi_{r}(\mu,m),\qquad\widetilde{\bm{L}}\psi_{r}(\mu,m)=\left({r}+{r}^{-1}\right)^{*}\psi_{r}(\mu,m),$ (5.6)
$\quad r=\exp\left[\frac{2\pi i}{N}(-ib\mu_{r}-m_{r})\right],\qquad\quad{r}^{*}=\exp\left[\frac{2\pi i}{N}(-ib^{-1}\mu_{r}+m_{r})\right].$ (5.7)
$\psi_{r}(\mu,m)$ is invariant manifestly under $(\mu_{r},m_{r})\to(-\mu_{r},-m_{r})$ and periodic under $m\to m+N$, $m_{r}\to m_{r}+N$. When $N=1$, $m=m_{r}=0$, $\psi_{r}$ reduces to Kashaev’s eigenfunctions [1] ($\phi(x,\lambda)$ in [7]) for the Dehn-twist operator.
@@
The following two lemmas state that the set of $\psi_{r}$ forms a complete distributional orthogonal basis for $\mathcal{H}$.
@@
The eigenstates $\psi_{r}$ satisfy the orthogonality
$\left\langle\psi_{r}\mid\psi_{r^{\prime}}\right\rangle=\sum_{m\in\mathbb{Z}/N\mathbb{Z}}\int\mathrm{d}\mu\,\psi_{r}(\mu,m)^{*}\psi_{r^{\prime}}(\mu,m)=\frac{1}{4}\rho(\mu_{r},m_{r})\delta_{r,r^{\prime}}.$ (5.8)
@@
$\rho(\mu_{r},m_{r})$ $=$ $N^{2}\left[\sin\left(\frac{2\pi}{N}(ib\mu_{r}+m_{r})\right)\sin\left(\frac{2\pi}{N}(-ib^{-1}\mu_{r}+m_{r})\right)\right]^{-1},$ (5.9)
$\delta_{r,r^{\prime}}$ $=$ $\delta(\mu_{r}-\mu_{r^{\prime}})\delta_{m_{r},m_{r^{\prime}}}+\delta(\mu_{r}+\mu_{r^{\prime}})\delta_{m_{r},-m_{r^{\prime}}}.$ (5.10)
@@
$(-1)^{m_{r}^{2}-m_{r^{\prime}}^{2}}e^{-\frac{i\pi(m_{r}^{2}-m_{r^{\prime}}^{2})}{N}+4i\pi x(\omega^{\prime\prime}-i\epsilon)}\frac{\gamma\left(-\lambda^{\prime}+x-\omega^{\prime\prime}+i\epsilon,m-{m_{r^{\prime}}}\right)\gamma\left(\lambda^{\prime}+x-\omega^{\prime\prime}+i\epsilon,m+{m_{r^{\prime}}}\right)}{\gamma\left(-\lambda+x+\omega^{\prime\prime}-i\epsilon,m-{m_{r}}\right)\gamma\left(\lambda+x+\omega^{\prime\prime}-i\epsilon,m+{m_{r}}\right)},$
where $\lambda^{\prime}={\mu_{r^{\prime}}}/{\sqrt{N}}$. We use the integration identity (A.18) to transform the ratio of two $\gamma$’s  [^5] The variables involved in (A.18) are given by $t=x$, $\alpha=\lambda^{\prime}-\omega^{\prime\prime}+i\epsilon,\ \beta=\lambda^{\prime}+\omega^{\prime\prime}-i\epsilon$, $d=m$, $a=m_{r}^{\prime}$, and $p=m_{r}$. We have $\mathrm{Im}\left(\alpha+\frac{c_{b}}{\sqrt{N}}\right)=\epsilon>0,\ \mathrm{Im}\left(\frac{c_{b}}{\sqrt{N}}-\beta\right)=\epsilon>0,\ \mathrm{Im}(\alpha-\beta)=2\epsilon-\frac{b+b^{-1}}{\sqrt{N}}<0$ for sufficiently small $\epsilon>0$. Therefore there exist an integration contour of $s$ with $\mathrm{Im}(\alpha-\beta)<\mathrm{Im}(s)<0$ allowing (A.18) to hold.
$\frac{\gamma\left(\lambda^{\prime}+x-\omega^{\prime\prime}+i\epsilon,m+{m_{r^{\prime}}}\right)}{\gamma\left(\lambda+x+\omega^{\prime\prime}-i\epsilon,m+{m_{r}}\right)}$ (5.12)
$=$ $\frac{\zeta_{0}}{\sqrt{N}}\sum_{c\in\mathbb{Z}/N\mathbb{Z}}\int\mathrm{d}s\frac{e^{\frac{2i\pi c(m+m_{r^{\prime}})}{N}+2i\pi s(x+\lambda^{\prime})}\gamma\left(-\lambda^{\prime}+\lambda+s+\omega^{\prime\prime}-2i\epsilon,-c-{m_{r^{\prime}}}+{m_{r}}\right)}{\gamma\left(s+\omega^{\prime\prime},-c\right)\gamma\left(-\lambda^{\prime}+\lambda+\omega^{\prime\prime}-2i\epsilon,{m_{r}}-{m_{r^{\prime}}}\right)},$